Optimized Model of Power Load Based on Grey Relation Analysis

Shijun Yang North China Electric Power University, China

Medium and long term load of power system has two layer trend of increscent and fluctuantat the same time. So the change of power load appeared very complicated nonlinear character. The optimized model of power load based on grey relation analysis increased the forecast precision from two ways. The one is improved the GM(1,1) model and founded the model storeroom. The other is chosen improved model through grey relation analysis.